sqrt(7x + 29) = x + 3
Please help, need all real and imaginary solutions

Answer by checkley77(12844)

sqrt(7x + 29) = x + 3 SQUARE BOTH SIDES
7X+29=X^2+6X+9
X^2+6X-7X+9-29=0
X^2-X-20=0
(X-5)(X+4)=0
X-5=0
X=5 ANS.
X+4=0
X=-4 ANS